OpenStudy (anonymous):

how would i graph y=x+2 i already did the first quadratic parabola which was y=x^2-2x+2

OpenStudy (shayaan_mustafa):

y=x+2 means graph is shifted two units up.

OpenStudy (anonymous):

 thats how my parobala looks like i just need help to solve y=x+2

OpenStudy (anonymous):

any specific way to show work ?

jhonyy9 (jhonyy9):

y=x+2 - will be a line

OpenStudy (anonymous):

i know but any way to show work?

jhonyy9 (jhonyy9):

what you wann again ?

OpenStudy (anonymous):

how do i show work for the line y=x+2

jhonyy9 (jhonyy9):

add values for x nd than calcule how many will be y

jhonyy9 (jhonyy9):

- like for example for x=0 than y=2 - for x=1 --- y=3

OpenStudy (anonymous):

still confuzed could you draw it please.

jhonyy9 (jhonyy9):